# Subtleties in the pseudomodes formalism

### v1

@@ II Computation of Effective Spectral Density
The form that the effective spectral density (16) of bath $\alpha$ represented by pseudomodes takes depends on whether the matrix $W_{\alpha}$ (15) is diagonal, diagonalizable, or non-diagonalizable. A pseudomode configuration is said to be diagonal, diagonalizable or non-diagonalizable when its matrix $W_{\alpha}$ has the corresponding property. These cases are described sequentially below. In general we might also have a matrix $W_{\alpha}$ which is block-diagonal, where some blocks are diagonalizable and some are not; in this case, each block can be treated separately and the resulting effective spectral density is merely the sum of the effective spectral densities from each block.
@@ II.1 Diagonal Case
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k}\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k}\zeta_{\alpha i,k}\zeta_{\alpha j,k}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\,.\end{cases}$ (17)
@@
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k}\frac{\mathfrak{Re}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (18)
@@
$J^{\rm eff}_{\alpha,ii}(\omega)=\sum_{k}\frac{\left|\zeta_{\alpha i,k}\right|^{2}\gamma_{\alpha k}}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,$, (19)
@@ II.2 Diagonalizable Case
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k}\kappa_{\alpha k,ij}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k}\kappa_{\alpha k,ij}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\end{cases}\,$, (23)
@@
$\kappa_{\alpha k,ij}=\sum_{pq}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{\alpha,pk}S_{\alpha,kq}^{-1}\,.$ (24)
Note that the diagonal elements $\kappa_{\alpha k,ii}$ always sum to a positive real number:
$\sum_{k}\kappa_{\alpha k,ii}>0\,.$ (25)
@@
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k}\frac{\mathfrak{Re}\left(\kappa_{\alpha k,ij}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\kappa_{\alpha k,ij}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (26)
Figure 2: (a) Comparison of fits to an effective spectral density from a diagonal configuration (sum of Lorentzians, with uncoupled pseudomodes) and from a diagonalizable one (sum of Lorentzians and Anti-Lorentzians, with pseudomodes coupled to one another). The true spectral density is $J(\omega)=\sqrt{1-\omega^{2}}$, which both fits model using 6 pseudomodes. The diagonal-case fit was obtained by brute-force optimization for the parameters $|\zeta_{k}|^{2},\varepsilon_{k},\gamma_{k}$, while for the diagonalizable case, Prony’s method (described in section III and appendix D) was used to obtain the best-fit $\kappa_{k},\varepsilon_{k},\gamma_{k}$. (b) Contributions to the effective spectral density in the diagonal case from each of the 6 pseudomodes; each mode contributes a Lorentzian term. (c) Contributions to the effective spectral density in the diagonalizable case from each of the 6 pseudomodes; each mode contributes the sum of a Lorentzian term and an Anti-Lorentzian term.
@@ II.3 Non-Diagonalizable Case
So long as $W_{\alpha}$ is diagonalizable, we can only end up with Lorentzian (20) and Anti-Lorentzian (21) terms in the spectral density matrix. However, in the case that $W_{\alpha}$ is non-diagonalizable (sometimes called “defective”), other functional forms can appear. An $n\times n$ matrix is non-diagonalizable if it has fewer than $n$ unique eigenvectors, and is “maximally defective” if it has only one eigenvector.
If $W_{\alpha}$ is non-diagonalizable, it can be written in Block Jordan form. Each eigenvalue of $W_{\alpha}$ corresponds to a $d_{i}\times d_{i}$ block, where $d_{i}$ is the algebraic multiplicity of that eigenvalue. For simplicity suppose that $W_{\alpha}$ has one maximally defective eigenvalue $w$; the general result then follows. In that case we can write
@@
The resulting contribution to the effective memory kernel (14) (for $t>0$) is then:
$\chi_{\alpha,ij}^{\rm eff}(t)=e^{wt}\sum_{p,q}\sum_{k=1}^{n}\sum_{l=1}^{k}\frac{1}{(k-l)!}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{pk}S^{-1}_{lq}t^{k-l}\,.$ (29)
We can see in particular that since $1\leq l\leq k\leq d$, the highest possible power of $t$ that we can get inside the summation is $t^{d-1}$. The general form of the memory kernel is thus
$\chi(t)\sim\sum_{k=0}^{d-1}\kappa_{k}t^{d}e^{wt}\,.$ (30)
@@
with system couplings $\zeta=\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T}$}. The effective memory kernel (14) then has an exponential component and a component that is $t$ times an exponential:
@@
$\delta t\left[|\zeta_{1}|^{2}-|\zeta_{2}|^{2}-i\left(\zeta_{1}\zeta_{2}^{*}+\zeta_{1}^{*}\zeta_{2}\right)\right]\Big\}\,.$ (34)
@@
If we take the case where $\eta=0$ and $\zeta_{2}=0$, then this $W$ represents two pseudomodes in a chain-like configuration, with the system coupled to mode $1$, which is then coupled to mode $2$, which is coupled to a residual bath, as shown in Fig. 3. In this case, the spectral density we obtain is a squared Lorentzian:
@@
Figure 3: The simplest non-diagonalizable pseudomode configuration, consisting of two pseudomodes. One mode is coupled to the system and the other is coupled to the residual bath. If the energies and couplings are chosen so that $W$ is non-diagonalizable, then the system feels an effective spectral density which is a squared Lorentzian, as in (36).
@@
These non-diagonalizable configurations allow for more variety in the types of effective spectral densities that can be obtained from the pseudomode method beyond only Lorentzians and Anti-Lorentzians, though the terms obtained this way can be closely approximated by just Lorentzians and Anti-Lorentzians. If we treat a non-diagonalizable configuration as the limit of a barely-diagonalizable configuration, then the resulting diagonalizable-type spectral density approaches a non-diagonalizable one, as shown in Appendix C. However, these non-diagonalizable configurations also provide clues into how to fit more general types of spectral densities. For example, from Table 1 we can see that a spectral density whose denominator contains terms of order up to $\omega^{6}$ might be able to be fit exactly by just 3 modes.
@@ III Few-Mode Fitting
    The couplings to the residual baths should be as small as possible, to justify the use of a local master equation approach to describe them.
  3. (iii)
@@
Items (ii) and (iii) exist somewhat in conflict, since if the $J_{\alpha,ij}$ are not sharply peaked then the fewer pseudomodes we use, the larger the couplings between those pseudomodes and the residual baths will be.
@@
Depending on whether we prioritize (ii) or (iii) more highly, two approaches exist. If we prioritize using a small number of modes, as is most common, we first perform a fit to the memory kernel and then try to find pseudomode parameters that give the correct amplitudes. This procedure is discussed in the remainder of this section. If instead we are willing to use a large number of pseudomodes to ensure the residual bath couplings will be small, one option is to proceed similarly to Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025) as described in section IV.
We will assume a diagonalizable configuration for this section where not otherwise specified. Thus, the effective memory kernel is given by (22) and the effective spectral density is given by (26).
III.1 Fitting the Memory Kernel with Prony’s Method
@@
We fit the memory kernel from each bath separately. We first fit each element of $\chi_{\alpha,ij}^{\rm eff}(t)$ as a sum of complex exponentials using Prony’s method Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022), an algorithm originally designed for discrete signal processing and detailed in Appendix D. Prony’s method works by fitting a discrete set of evenly spaced points with complex exponentials; we found the best results by performing optimization of the Prony’s method fit for a variable set of discrete points. This gives us the set of $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ from which to construct the effective memory kernel (22). Since the exponents in $\chi_{\alpha,ij}^{\rm eff}$ are independent of $i,j$, in the end each $i,j$ element of the memory kernel matrix should be fit by a sum of exponentials with the same exponents but different amplitudes $\kappa_{\alpha k,ij}$. Caution must be used here, since the fit obtained may produce a spectral density that is not positive everywhere or even has non-decaying exponentials. Few-mode fitting has also been performed by brute-force optimization over the pseudomode parameters without first fitting the memory kernel Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); doing so guarantees a fit with physical pseudomode parameters, but is much less efficient computationally. Note, also, that some spectral densities will be easier to fit than others. In particular, if we have a wideband $J(\omega)=\Gamma_{0}$ for $\omega_{m}\leq\omega\leq\omega_{M}$ and 0 otherwise, then the corresponding memory kernel is
@@ III.2 Obtaining Pseudomode Parameters
Choosing the parameters $\Lambda,\Gamma,\zeta$ to obtain the desired amplitudes $\kappa$ is an Inversion Problem, and it remains the greatest challenge in implementing the pseudomode technique. The relation (24) between the constants $\kappa_{\alpha k,ij}$ and the actual pseudomode parameters $\Lambda_{\alpha},\Gamma_{\alpha},\zeta_{\alpha}$ is highly nonlinear, and there are many more degrees of freedom in these parameters than are required. The fit with Prony’s method for $n$ pseudomodes gives us $4n$ real parameters: $n$ complex amplitudes and $n$ complex exponents. For the actual pseudomode parameters, we have $n^{2}+3n$ real free parameters: $n^{2}$ from $\Lambda$ (Hermitian), $n$ from $\Gamma$ (real and diagonal), and $2n$ from $\zeta$ (complex). However, not all of these free parameters actually contribute due to the structure of the $\kappa$’s (24), and there are a number of constraints that are not readily apparent. For example, a simple one is that, because we always see pairs $\zeta_{\alpha i}^{*}\zeta_{\alpha j}$, the vector $\zeta_{\alpha i}$ is invariant under a global phase $\zeta\to e^{i\phi}\zeta$.
@@
where $M$ is a diagonal matrix whose entries are the eigenvalues of $W$ obtained from the Prony’s method fit. While the eigenvalues of $W$ are specified, $S$ is not. Note that if we transform $W$ by some unitary transformation, then as long as we transform $\zeta$ by the same unitary we will obtain the same effective memory kernel (22):
@@
We can write any solution to (44) by first fixing the values of $u_{j}$ and then taking
@@
where $B_{\perp}$ can be any positive matrix that satisfies $B_{\perp}u=0$ and makes (46) Hermitian.
@@
This approach can always be used to obtain a set of pseudomode parameters whose effective spectral density will exactly match the effective spectral density obtained from fitting using Prony’s method, provided that Prony’s method outputs a set of $\kappa_{k}$ that satisfy (25). However, this approach is not guaranteed to give a matrix $\Gamma>0$.
It is worth emphasizing again here the sheer amount of freedom involved in this process. In (45) we can choose anything for the $u_{j}$. Then in (46), $B_{\perp}$ can by any positive-definite $n\times n$ matrix satisfying $B_{\perp}u=0$; alternately, we can choose any positive block $B$ in (49) and any value of $A_{12}$ satisfying (51). A naturally interesting question is then: for what fit parameters $\kappa_{k},\varepsilon_{k},\gamma_{k}$ will there be a choice of $u_{j},B_{\perp}$ such that we obtain $\Gamma>0$? We were unable to find a general answer to this question, but we explore it further for the simplest case of 2 pseudomodes in Appendix B. For that case, we show that it is possible to have $\Gamma>0$ if and only if the effective spectral density from the fit parameters is positive everywhere.

### v2

@@ II Computation of Effective Spectral Density
The form that the effective spectral density (17) of bath $\alpha$ represented by $n_{\alpha}$ pseudomodes takes depends on whether the matrix $W_{\alpha}$ (16) is diagonal, diagonalizable, or non-diagonalizable. A pseudomode configuration is said to be diagonal, diagonalizable or non-diagonalizable when its matrix $W_{\alpha}$ has the corresponding property. These cases are described sequentially below. In general we might also have a matrix $W_{\alpha}$ which is block-diagonal, where some blocks are diagonalizable and some are not; in this case, each block can be treated separately and the resulting effective spectral density is merely the sum of the effective spectral densities from each block.
While all the results shown in this section follow straightforwardly from (15) and (17), we show them explicitly, as we will use these results for the remainder of the paper. Writing these results explicitly also makes clear the difference in the effective spectral densities that can be obtained using coupled modes as opposed to uncoupled modes, and why it is often desirable to use coupled modes. Additionally, the case where $W_{\alpha}$ is not diagonalizable has not, to the best of our knowledge, been previously explored.
@@ II.1 Diagonal Case
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k=1}^{n_{\alpha}}\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k=1}^{n_{\alpha}}\zeta_{\alpha i,k}\zeta_{\alpha j,k}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\,.\end{cases}$ (18)
@@
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k=1}^{n_{\alpha}}\frac{\mathfrak{Re}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (19)
@@
$J^{\rm eff}_{\alpha,ii}(\omega)=\sum_{k=1}^{n_{\alpha}}\frac{\left|\zeta_{\alpha i,k}\right|^{2}\gamma_{\alpha k}}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,$, (20)
@@
Figure 2: (a) Comparison of fits to an effective spectral density from a diagonal configuration (sum of Lorentzians, with uncoupled pseudomodes) and from a diagonalizable one (sum of Lorentzians and Anti-Lorentzians, with pseudomodes coupled to one another). The true spectral density is $J(\omega)=\sqrt{1-\omega^{2}}$, which both fits model using 6 pseudomodes. The diagonal-case fit was obtained by brute-force optimization for the parameters $|\zeta_{k}|^{2},\varepsilon_{k},\gamma_{k}$, while for the diagonalizable case, Prony’s method (described in section III and appendix B) was used to obtain the best-fit $\kappa_{k},\varepsilon_{k},\gamma_{k}$. (b) Contributions to the effective spectral density in the diagonal case from each of the 6 pseudomodes; each mode contributes a Lorentzian term. (c) Contributions to the effective spectral density in the diagonalizable case from each of the 6 pseudomodes; each mode contributes the sum of a Lorentzian term and an Anti-Lorentzian term. The parameters from optimization used in these fits are given in Appendix F.
@@ II.2 Diagonalizable Case
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k=1}^{n_{\alpha}}\kappa_{\alpha k,ij}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k=1}^{n_{\alpha}}\kappa_{\alpha k,ij}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\end{cases}\,$, (24)
@@
$\kappa_{\alpha k,ij}=\sum_{p,q=1}^{n_{\alpha}}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{\alpha,pk}S_{\alpha,kq}^{-1}\,.$ (25)
Note that the diagonal elements $\kappa_{\alpha k,ii}$ may be complex, but always sum to a positive real number:
$\sum_{k=1}^{n_{\alpha}}\kappa_{\alpha k,ii}>0\,.$ (26)
This fact will become a key constraint when choosing pseudomode parameters to fit a spectral density.
@@
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k=1}^{n_{\alpha}}\frac{\mathfrak{Re}\left(\kappa_{\alpha k,ij}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\kappa_{\alpha k,ij}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (27)
@@ II.3 Non-Diagonalizable Case
So long as $W_{\alpha}$ is diagonalizable, we can only end up with Lorentzian (21) and Anti-Lorentzian (22) terms in the spectral density matrix. However, in the case that $W_{\alpha}$ is non-diagonalizable (sometimes called “defective”), which in the language of non-Hermitian Hamiltonians corresponds to an exceptional point in the pseudomodes’ non-Hermitian Hamiltonian $iW$, other functional forms can appear.
An $n\times n$ matrix is non-diagonalizable if it has fewer than $n$ unique eigenvectors, and is “maximally defective” if it has only one eigenvector. If $W_{\alpha}$ is non-diagonalizable, it can be written in Block Jordan form. Each eigenvalue of $W_{\alpha}$ corresponds to a $d_{i}\times d_{i}$ block, where $d_{i}$ is the algebraic multiplicity of that eigenvalue. For simplicity suppose that $W_{\alpha}$ has one maximally defective eigenvalue $w$; the general result then follows. In that case we can write
@@
The resulting contribution to the effective memory kernel (15) (for $t\geq 0$) is then:
$\chi_{\alpha,ij}^{\rm eff}(t)=e^{wt}\sum_{p,q=1}^{n_{\alpha}}\sum_{k=1}^{n_{\alpha}}\sum_{l=1}^{k}\frac{1}{(k-l)!}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{pk}S^{-1}_{lq}t^{k-l}\,.$ (30)
We can see in particular that since $1\leq l\leq k\leq d$, the highest possible power of $t$ that can appear inside the summation is $t^{d-1}$. The general form of the memory kernel for $t\geq 0$ is thus
$\chi(t)\sim\sum_{k=0}^{d-1}\kappa_{k}t^{k}e^{wt}\,.$ (31)
@@
with system couplings $\zeta=\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T}$}. This corresponds to the configuration shown in Fig. 3(a). The effective memory kernel (15) then has an exponential component and a component that is $t$ times an exponential:
@@
$\delta t\left[|\zeta_{1}|^{2}-|\zeta_{2}|^{2}-i\left(\zeta_{1}\zeta_{2}^{*}+\zeta_{1}^{*}\zeta_{2}\right)\right]\Big\}\,.$
@@
If we take the case where $\eta=0$ and $\zeta_{2}=0$, then this $W$ represents two pseudomodes in a chain-like configuration, with the system coupled to mode $1$, which is then coupled to mode $2$, which is coupled to a residual bath, as shown in Fig. 3(b). In this case, the spectral density we obtain is a squared Lorentzian:
@@
Figure 3: (a) The most general non-diagonalizable pseudomode configuration, consisting of two pseudomodes with $W$ given by (33). (b) The simplest non-diagonalizable pseudomode configuration, with $\zeta_{2}=\eta=0$. One mode is coupled to the system and the other is coupled to the residual bath. If the energies and couplings are chosen so that $W$ is non-diagonalizable, then the system feels an effective spectral density which is a squared Lorentzian, as in (36).
@@
These non-diagonalizable configurations allow for more variety in the types of effective spectral densities that can be obtained from the pseudomode method beyond only Lorentzians and Anti-Lorentzians, though the terms obtained this way can be closely approximated by just Lorentzians and Anti-Lorentzians. If we treat a non-diagonalizable configuration as the limit of a barely-diagonalizable configuration, then the resulting diagonalizable-type spectral density approaches a non-diagonalizable one, as shown in Appendix C. However, these non-diagonalizable configurations also provide clues into how to fit more general types of spectral densities. For example, from Table 1 we can see that a spectral density whose denominator contains terms of order up to $\omega^{6}$ might be able to be fit exactly by just 3 modes. An additional example of a non-diagonalizable configuration is shown in Appendix D.
@@ III Few-Mode Fitting
Depending on how strongly we prioritize (ii), multiple approaches exist. While most approaches prioritize using a small number of modes, if instead we are willing to use a large number of pseudomodes, one option is to proceed similarly to Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025) as described in section IV. In the remainder of this section, we focus on few-mode fitting. We will assume a diagonalizable configuration for this section where not otherwise specified. Thus, the effective memory kernel is given by (24) and the effective spectral density is given by (27).
The most common approach to few-mode fitting is to first fit the memory kernel in terms of the parameters $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$, and then to try to choose pseudomode parameters $\Lambda,\Gamma,\zeta$ that give rise to the same effective memory kernel. We focus on this method. One might alternately choose to fit the correlation functions (7) (or their Fourier transforms), thus encoding the temperatures and chemical potentials of the original baths into the pseudomode parameters, rather than the residual baths Pleasance et al. (2020); Menczel et al. (2024). Few-mode fitting has also been performed by brute- force optimization over the pseudomode parameters without first fitting the memory kernel Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); doing so guarantees a fit with physical pseudomode parameters (which we shall see is not always guaranteed otherwise), but is much less efficient computationally. A third method was recently proposed in Ref. Huang et al. (2025), which instead maps (17) onto the so-called control-theoretic realization problem Mayo and Antoulas (2007); Antoulas and Anderson (1986).
In the remainder of this section, we present a new method for choosing the pseudomode parameters, for which the only optimization step is the fitting of the memory kernel. Using this method, the pseudomode parameters can be chosen to exactly match the $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ obtained by fitting the memory kernel.
III.1 Fitting the Memory Kernel
@@
We fit the memory kernel from each bath separately. We first fit each element of $\chi_{\alpha,ij}^{\rm eff}(t)$ as a sum of complex exponentials using techniques from discrete signal processing, such as Prony’s method de Prony (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) or the ESPRIT algorithm Roy et al. (1986); Paulraj et al. (1986); Sahnoun et al. (2017) (see Appendix B). These methods work by sampling the true $\chi(t)$ at discrete, evenly spaced times, and then fitting those discrete points with complex exponentials. We found the best results by using ESPRIT, as it is more stable with respect to the set of points at which $\chi$ is sampled than Prony’s method. This gives us the set of $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ from which to construct the effective memory kernel (24). Since the exponents in $\chi_{\alpha,ij}^{\rm eff}$ are independent of $i,j$, in the end each $i,j$ element of the memory kernel matrix should be fit by a sum of exponentials with the same exponents but different amplitudes $\kappa_{\alpha k,ij}$. Caution must be used here, since the fit obtained may produce a spectral density that is not positive everywhere or even has non-decaying exponentials. Note, also, that some spectral densities will be easier to fit than others. In particular, if we have a wideband $J(\omega)=\Gamma_{0}$ for $\omega_{m}\leq\omega\leq\omega_{M}$ and 0 otherwise, then the corresponding memory kernel is
@@ III.2 Obtaining Pseudomode Parameters
Choosing the parameters $\Lambda,\Gamma,\zeta$ to obtain the desired amplitudes $\kappa$ is an Inversion Problem, and it remains the greatest challenge in implementing the pseudomode technique. The relation (25) between the constants $\kappa_{\alpha k,ij}$ and the actual pseudomode parameters $\Lambda_{\alpha},\Gamma_{\alpha},\zeta_{\alpha}$ is highly nonlinear, and there are many more degrees of freedom in these parameters than are required. The fit with Prony’s method (or similar algorithms) for $n$ pseudomodes gives us $4n$ real parameters: $n$ complex amplitudes and $n$ complex exponents. For the actual pseudomode parameters, we have $n^{2}+3n$ real free parameters: $n^{2}$ from $\Lambda$ (Hermitian), $n$ from $\Gamma$ (real and diagonal), and $2n$ from $\zeta$ (complex). However, not all of these free parameters actually contribute due to the structure of the $\kappa$’s (25), and there are a number of constraints that are not readily apparent. For example, a simple one is that, because we always see pairs $\zeta_{\alpha i}^{*}\zeta_{\alpha j}$, the vector $\zeta_{\alpha i}$ is invariant under a global phase $\zeta\to e^{i\phi}\zeta$. Similarly, we can express the pseudomode operators $a_{\alpha i}$ in any basis without changing the dynamics. We choose the basis in which $\Gamma$ is diagonal so that the resulting master equation will have local dissipators. Physically, this is equivalent to the specification in (I.2) and (14) that each pseudomode couples to its own residual bath and that each of those residual baths has a constant spectral density $\Gamma_{\alpha,kk}$.
@@
where $M$ is a diagonal matrix whose entries are the eigenvalues of $W$ obtained from the fitting the memory kernel. While the eigenvalues of $W$ are specified, $S$ is not. Note that if we transform $W$ by some unitary transformation, then as long as we transform $\zeta$ by the same unitary we will obtain the same effective memory kernel (24):
@@
We can write any solution to (43) by first fixing the values of $u_{j}$, which must be nonzero, and then taking
@@
where $B_{\perp}$ can be any positive matrix that satisfies $B_{\perp}u=0$ and makes (45) Hermitian. The choice of $B_{\perp}$ matrix will not affect the effective spectral density, but will affect the pseudomode parameters $\Lambda,\Gamma$. In particular, this will affect the rates in the master equation, which are given by the diagonal entries of $\Gamma$.
@@
This approach can always be used to obtain a set of pseudomode parameters whose effective spectral density will exactly match the effective spectral density obtained from fitting the memory kernel, provided that the fitting procedure used outputs a set of $\kappa_{k}$ that satisfy (26). However, this approach is not guaranteed to give a matrix $\Gamma>0$.
Figure 4: Fits to (a) a memory kernel of the form (57) and (b) the corresponding spectral density (56), with $\Delta=0.04$ and $\delta=6.0$. The blue curves are obtained by fitting (57) with the ESPRIT algorithm (see Appendix B). After following the inversion procedure of section III.2, optimizing over the vector $u$ (44) and matrix $B$ (48) to minimize any negative elements of $\Gamma$, one obtains the pseudomode parameters $\Lambda,\zeta,\Gamma$ corresponding to that fit. Finally, setting any negative elements of the matrix $\Gamma$ (which would cause negative rates in the corresponding master equation) gives rise to the modified fit shown in red. These fits are compared with the best fit using uncoupled modes, which was obtained using brute-force optimization. The pseudomode parameters appearing in these fits are given in Appendix F.
It is worth emphasizing again here the sheer amount of freedom involved in this process. In (44) we can choose anything for the $u_{j}$. Then in (45), $B_{\perp}$ can by any positive-definite $n\times n$ matrix satisfying $B_{\perp}u=0$; alternately, we can choose any positive block $B$ in (48) and any value of $A_{12}$ satisfying (50). A naturally interesting question is then: for what fit parameters $\kappa_{k},\varepsilon_{k},\gamma_{k}$ will there be a choice of $u_{j},B_{\perp}$ such that we obtain $\Gamma>0$? We were unable to find a general answer to this question, but we explore it further for the simplest case of 2 pseudomodes in Appendix E. For that case, we show that it is possible to have $\Gamma>0$ if and only if the effective spectral density from the fit parameters is positive everywhere.
III.3 Numerical Implementation
As an example, we now apply this inversion procedure to a spectral density which is the sum of a semicircular term with a wide Lorentzian:
$J(\omega)=\sqrt{(1-\omega)^{2}}\,+\frac{\Delta\delta}{\omega^{2}+(\delta/2)^{2}}\,.$ (56)
We choose this spectral density because fermionic system coupled to an infinite, homogenous wire will have a semi-elliptical spectral density Ferracin et al. (2024). We add the wide Lorentzian to represent how no physical system has true hard cutoff frequencies in its spectral density, and because the methods of section III.1 struggle to fit the sharp corners of a semicircular spectral density without producing an effective spectral density that is negative somewhere.
The corresponding memory kernel is
$\chi(t)=\frac{1}{2t}\mathcal{J}_{1}(t)+\Delta e^{-\delta t}$ (57)
where $\mathcal{J}_{1}(t)$ is a Bessel function. We first fit $\chi(t)$ with pseudomodes using the ESPRIT algorithm Paulraj et al. (1986); Roy et al. (1986); Sahnoun et al. (2017) (see Appendix B). Then we optimize over the values of the vector $u$ from (44) and the matrix $B$ from (48), trying first to make any negative rates as small as possible, and then minimizing the largest rate if $\Gamma>0$.
As an example, we construct a 6-mode fit to the spectral density (56), using $\Delta=0.04$ and $\delta=6.0$. When optimizing over $u,B$, we obtain five positive rates and one negative rate. The spectral density and memory kernel and their corresponding fits are shown in figure 4 (a) and (b) respectively, along with the effective fits that we obtain if we set the one negative rate to zero. We compare our results with an optimized fit involving only uncoupled modes. While in this case we are unable to obtain a fit with only nonnegative rates, our modified fit where the negative rates are discarded still fits the true memory kernel quite well at both short and long times with only minor deviations at the peaks and troughs of $\chi(t)$ at intermediate times.